A right-circular cone is inscribed in a sphere having a fixed radius of 10 in.
Express the volume of the cone as a function of its radius.
1
Expert's answer
2020-09-15T15:22:51-0400

By Pythagorean theorem we have 102=R2+(H-10)2 (see the picture).

Therefore, H = 10+(100-R2)1/2; and V = "\pi" (10+(100-R2)1/2)R2/3.